U2 - 6

Your firm has identified three potential investment projects. The projects and their cash flows are shown here: Cash Flow Today (millions) -$14 $4 S22 Cash Flow in One Year (millions) 523 $4 - $7 Project Suppose all cash flows are certain and the risk-free interest rate is 9% a. What is the NPV of each project? b. If the firm can choose only one of these projects, which should it choose? c. If the firm can choose any two of these projects, which should it choose?

?(Round to two decimal? places.)
Answer #1

NPV is calculated by discounting the cashflows

PV = C/(1+r)^n

C - Cashflow

r - Discount rate

n - years to the cashflow

Project A:

NPV = -14 + 23/(1+0.09)^1 = 7.10 million

Project B:

NPV = 4 + 4/(1+0.09)^1 = 7.67 million

Project C:

NPV = 22 - 7 /(1+0.09)^1 = 15.58 million

b.

Choose project C, since it has the highest NPV.

c.

Choose projects C and B, since they have higher NPV compared to project A.